OpenStudy (anonymous):

Need help, please. A segment of a circle has a 120 degree arc and a chord of 8 square root 3 in. Find the area of the segment.

OpenStudy (anonymous):

if we calculate the area of the circle, divide it by 3 to get the 120 degree slice and then subtract the area of the isosceles triangle there, we'll have the area of the segment: area of circle = pi*r^2 = pi*8*8 = 64 pi area of slice = 64 pi / 3 area of triangle = h * chord / 2 = 4 * 8sqrt(3) / 2 = 16* sqrt(3) area of segment = 64 pi / 3 - 16*sqrt(3)

OpenStudy (anonymous):

Thank you Heather, that was correct. I'll try some on my own and see if I can do this now.

OpenStudy (anonymous):

No problem :)